I am having trouble with a problem, I think I found the answer but, I am not sure. My question is Solve the subtraction problem 8/9 - 3/4? I think the answer is 5/36? Am I right, Please let me know.
Answer by Earlsdon(6294)

Congratulations! You are absolutely correct.
8/9 - 3/4 The common denominator is 36. Writing the equivalent fractions:
8/9 = 32/36
3/4 = 27/36
Now you have: 32/36 - 27/36 = 5/36